Let z=z1/z2 where z1=tantheta +i and z2=z1. Find an expression for zn with n E N
1
Expert's answer
2021-07-19T05:44:03-0400

"Z_0=0\\
Z_1=i+1\\
Z_2=i+1\\
Z_3=\bar{Z_2}+(-1)^2+1=3i+3\\
Z_4=\bar{Z_3}+(-1)^3=4-4i\\
So, Z_n = (-1)^{n+1}*ni+n"